A waste disposal company is considering the replacement of one of its aging trucks. The key parameters of the three trucks under scrutiny are provided below.

Parameters

Delta

Epsilon

Zeta

1. Initial Cost

($)

250,000

375,000

450,000

2. Revenues

($)

230,000 at EOY1

increasing by 2.5% annually thereafter

195,000 at EOY1

increasing by 3,000 annually thereafter

235,000 at EOY1

decreasing by 1% annually thereafter

3. Operating

costs ($)

140,000 at EOY1

decreasing by 2,000 annually thereafter

125,000 at EOY1

decreasing by 2% annually thereafter

EOY1-EOY4 = 125,000;

EOY5-EOY8 = 135,000

EOY9-EOY12 = 170,000

EOY13-EOY16 = 190,000

4. End-of-life salvage value ($)

-20,000

7,000

-20,000

5. Useful life

(years)

4

8

16

  • EOY = End-of-year
  • Industry Standard = 4 years
  • MARR = 10%
  1. Delta’s Net Future Worth (NFW) (rounded to the nearest $100) of
    a) $78,800; b) $81,900; c) $82,900; d) $84,300.
  1. Epsilon’s Net Present Worth (NPW) (rounded to the nearest $100) is
    a) $89,000; b) $93,700; c) $94,100; d) $95,200.
  1. Zeta’s Net Future Worth (NFW) (rounded to the nearest $100) after 16 years is

a) $678,900; b) $698,600; c) $709,100; d) $743,300.

  1. Delta’s Annual Equivalent Worth (AEW) (rounded to the nearest $100) is
    a) $17,700; b) $18,500; c) $18,900; d) $19,200.
  1. Zeta’s AEW (rounded to the nearest $100) over 30 years (it was repeated several times) is

a) $19,200; b) $19,700; c) $20,100; d) $21,800.

  1. The best truck based on the NFW method is
    a) Delta; b) Epsilon; c) Zeta.
  1. The best truck based on the AEW method is
    a) Delta; b) Epsilon; c) Zeta.
  1. Based on the simple payback method, Delta’s recovery period (to the nearest half or full year) is

a) 2.5; b) 3.0; c) 3.5; d) 4.0.

  1. Based on the simple payback method, Epsilon’s “project balance” after 2 years (rounded to the nearest $100) is

a) $-229,500; b) $-155,900; c) $-59,100; d) $98,800.

  1. Based on the discounted payback method, Epsilon’s recovery period (to the nearest half or full year) is

a) 5.0; b) 6.0; c) 6.5; d) 7.5.

.

  1. Based on the discounted payback method, Epsilon’s “project balance” after 3 years (rounded to the nearest $100) is

a) $-296,300; b) $-250,400; c) $123,400; d) $129,800.

  1. Based on the simple payback method, Zeta’s recovery period (rounded to the nearest half or full year) is

a) 3.5; b) 4.0; c) 4.5; d) 5.5.

  1. The best truck based on the simple payback method is
    a) Delta; b) Epsilon; c) Zeta.
  1. Delta’s benefit/cost (B/C) ratio (second decimal; no rounding) is
    a) 0.98; b) 1.03; c) 1.08; d) 1.10.
  1. Zeta’s benefit/cost (B/C) ratio (second decimal; no rounding) is
    a) 0.99; b) 1.02; c) 1.06; d) 1.09.
  1. The incremental B/C ratio (second decimal; no rounding) between the Delta and Epsilon trucks is

a) 0.90; b) 0.99; c) 1.03; d) 1.08.

  1. The incremental B/C ratio (second decimal; no rounding) between the Epsilon and Zeta trucks is

a) 0.98; b) 1.05; c) 1.07; d) 1.11.

  1. Delta’s Internal Rate of Return (IRR) (first decimal; no rounding) is
    a) 18.1%; b) 19.8%; c) 20.8%; d) 21.1%.
  1. Epsilon’s Internal Rate of Return (IRR) (first decimal; no rounding) is
    a) 16.0%; b) 16.8%; c) 17.0%; d) 17.3%.
Answer #1

For each truck,

Annual net benefit (NAB) = Annual Revenue - Annual Costs

NAB in the last year will be higher by the value of respective salvage value.

NPW of Epsilon and NFW for Delta and Zeta are computed as follows.

EPSILON
EOY (N) Revenue ($) Cost ($) NAB ($) PVIF(10%, N) Discounted NAB ($)
(A) (B) (C)=(A)-(B) (D) (C) x (D)
0 3,75,000 -3,75,000 1.0000 -3,75,000
1 1,95,000 1,25,000 70,000 0.9091 63,636
2 1,98,000 1,22,500 75,500 0.8264 62,397
3 2,01,000 1,20,050 80,950 0.7513 60,819
4 2,04,000 1,17,649 86,351 0.6830 58,979
5 2,07,000 1,15,296 91,704 0.6209 56,941
6 2,10,000 1,12,990 97,010 0.5645 54,760
7 2,13,000 1,10,730 1,02,270 0.5132 52,481
8 2,23,000 1,08,516 1,14,484 0.4665 53,408
NPW ($) = 88,420
DELTA
EOY (N) Revenue ($) Cost ($) NAB ($) FVIF(10%, N) Discounted NAB ($)
(A) (B) (C)=(A)-(B) (D) (C) x (D)
0 2,50,000 -2,50,000 1.4641 -3,66,025
1 2,30,000 1,40,000 90,000 1.3310 1,19,790
2 2,35,750 1,38,000 97,750 1.2100 1,18,278
3 2,41,644 1,36,000 1,05,644 1.1000 1,16,208
4 2,27,685 1,34,000 93,685 1.0000 93,685
NFW ($) = 81,935
ZETA
EOY (N) Revenue ($) Cost ($) NAB ($) FVIF(10%, N) Compounded NAB ($)
(A) (B) (C)=(A)-(B) (D) (C) x (D)
0 4,50,000 -4,50,000 4.5950 -20,67,738
1 2,35,000 1,25,000 1,10,000 4.1772 4,59,497
2 2,32,650 1,25,000 1,07,650 3.7975 4,08,801
3 2,30,324 1,25,000 1,05,324 3.4523 3,63,605
4 2,28,020 1,25,000 1,03,020 3.1384 3,23,322
5 2,25,740 1,35,000 90,740 2.8531 2,58,892
6 2,23,483 1,35,000 88,483 2.5937 2,29,501
7 2,21,248 1,35,000 86,248 2.3579 2,03,368
8 2,19,035 1,35,000 84,035 2.1436 1,80,137
9 2,16,845 1,70,000 46,845 1.9487 91,288
10 2,14,677 1,70,000 44,677 1.7716 79,147
11 2,12,530 1,70,000 42,530 1.6105 68,495
12 2,10,404 1,70,000 40,404 1.4641 59,156
13 2,08,300 1,90,000 18,300 1.3310 24,358
14 2,06,217 1,90,000 16,217 1.2100 19,623
15 2,04,155 1,90,000 14,155 1.1000 15,571
16 1,82,114 1,90,000 -7,886 1.0000 -7,886
NFW ($) = 7,09,137

Therefore,

(Question 1) Option (b)

(Question 2) Option (a)

(Question 3) Option (c)

(Question 4) Option (a)

AEW for Delta = NFW / F/A(10%, 4) = 81,935 / 4.641 = 17,655